OpenStudy (anonymous):

find dy/dx using implicit differentiation. compare your answer with the result obtained by first solving for y as a function of x then taking the derivative. e^xy-y=4

OpenStudy (anonymous):

\[e ^{xy-y} \] ?

OpenStudy (anonymous):

its written as e^x y-y=4

OpenStudy (anonymous):

its asking for implicit differentiation dy/dx= ?

OpenStudy (anonymous):

as well as solving for y first then take the derivative, so y= ? and dy/dy=?

OpenStudy (anonymous):

can't help you until I know what the problem is..

OpenStudy (anonymous):

that is the problem the equation is e^x y-y=4

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

I'll assume it's \[e ^{x} (y-y) = 4\]

OpenStudy (anonymous):

e^x*0 =4 0=4 false.

OpenStudy (anonymous):

IS IT \[e ^{xy} - y =4\] OR \[e ^{xy- y} =4\]

OpenStudy (anonymous):

it just says e^xy-y=4 that is exactly how it is written

OpenStudy (anonymous):

cool.

OpenStudy (anonymous):

thanks? no need to be rude here..

OpenStudy (anonymous):

Yes. Excuse me for bullying you on the internet. Which version of the problem would you prefer to do?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

\[e ^{xy} -y =4\] or \[e ^{xy -y} =4\]

OpenStudy (anonymous):

is it not possible to solve a problem given e^xy-y=4, because thats how it is written online, is that incorrect or something?

OpenStudy (anonymous):

It's entirely possible to solve it either way. Which one shall we solve?

OpenStudy (anonymous):

lets go with the first one please

OpenStudy (anonymous):

that tyou wrote

OpenStudy (anonymous):

Excellent choice. (e^u) ' = (e^u)*du (e^(xy)) ' = (e^xy)*(xy' +y)

OpenStudy (anonymous):

(e^xy)*(xy' +y) - y' =0

OpenStudy (anonymous):

group terms with y' on one side of the equation, factor out y', divide both sides by the factor on y'

OpenStudy (anonymous):

so I can subtract thewhole term e^xy from the other side?

OpenStudy (anonymous):

I don't know what you mean. You can do anything that the rules of algebra allow.

OpenStudy (anonymous):

is this fine? (xy'+y)-y'=-e^xy

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

I thought we had to isolate the y' so so far I moved over the e^xy term

OpenStudy (anonymous):

4*2 = 8 -4 -4 2 = 4

OpenStudy (anonymous):

seem reasonable?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

(e^xy)(xy'+y)=y'

OpenStudy (anonymous):

?

OpenStudy (anonymous):

group terms with y' on one side of the equation

OpenStudy (anonymous):

e^xy=-xy'+y-y' ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

can you please just help me..

OpenStudy (anonymous):

Perhaps it is \(e^xy-y = 4\)?

OpenStudy (anonymous):

thats the problem how it is exactly written

OpenStudy (anonymous):

I need to find the implicit differentiation of it, do you know how to do that?

OpenStudy (anonymous):

rolypoly

OpenStudy (anonymous):

Do you know who to differentiate \(e^xy\)?

OpenStudy (anonymous):

*how

OpenStudy (anonymous):

oh not for this problem, e^x would just be e^x but theres nothing to take the derivative of the y's..

OpenStudy (anonymous):

You need to apply product rule when differentiating that guy. Perhaps you would like to try again.

OpenStudy (anonymous):

wouldnt it just be e^xy dy/dx?

OpenStudy (anonymous):

Not really. Product rule: uv' + u'v

OpenStudy (anonymous):

In this case, u = e^x , v=y v' = y' Can you try again?

OpenStudy (anonymous):

e^x dy/dx+e^x(y)

OpenStudy (anonymous):

Yup. \[e^xy-y=4\] Diff. both sides w.r.t. x What do you get?

OpenStudy (anonymous):

e^x dy/dx+e^x(y)-y=4 what does w.r.t. x mean?

OpenStudy (anonymous):

so do i have to get the e^x dy/dx separate?

OpenStudy (anonymous):

w.r.t. = with respect to e^x dy/dx+e^x(y) <- right -y=4 <- not right.

OpenStudy (anonymous):

And not yet, you can't separate them at the moment, until you get this step correct.

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so i caan subtract y from each side first

OpenStudy (anonymous):

No! Differentiate both sides with respect to x means that you have to differentiate EVERY TERMS on both sides with respect to x.

OpenStudy (anonymous):

would it be e^x dy/dx+e^x(y)-y dy/dx=4?

OpenStudy (anonymous):

No, -y dy/dx and are not right.

OpenStudy (anonymous):

dy/dx = dy/dx, nothing more d/dx (constant) = 0, as always.

OpenStudy (anonymous):

oh ok sorry the y does not get dy/dx bc it is not being differntiated so its e^x dy/dx+e^x(y)=0

OpenStudy (anonymous):

No... d/dx (y) = dy/dx = y' Similarly, d/dx (p) = dp/dx = p'

OpenStudy (anonymous):

uhh im confused :/

OpenStudy (anonymous):

can you please guide me im confused and i dont have much time left for this question :/

OpenStudy (anonymous):

For the question find dy/dx of the equation y = 2x, how do you find it?

OpenStudy (anonymous):

2x*x ?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

i think for my actualy problem is it y/e^xy

OpenStudy (anonymous):

y=e^x

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

what am i doing wrong :/

OpenStudy (anonymous):

How do you get y=e^x? What is y/e^xy?

OpenStudy (anonymous):

i dont know im just taking a guess..

OpenStudy (anonymous):

my times running out :/

OpenStudy (anonymous):

e^xy(xy)'-y'=0 ?

OpenStudy (anonymous):

Almost correct! For the part e^xy(xy)', it's not correct, but the -y'=0 part is correct.

OpenStudy (anonymous):

e^x dy/dx+e^x(y)-y'=0

OpenStudy (anonymous):

Yes. That's correct. Now, dy/dx can be rewrite as y' Make y' as the subject, can you do it?

OpenStudy (anonymous):

so you mean it looks like e^x y'+e^x(y)-y'=0

OpenStudy (anonymous):

and isoloate the y'

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

so can i subtract the e^xy'+e^x(y) all of that?

OpenStudy (anonymous):

Subtract? why?

OpenStudy (anonymous):

to the other side to get y' by itself

OpenStudy (anonymous):

e^x y'+e^x(y)-y'=0 Rearrange it. e^x y'-y'+e^x(y)=0 Subtract e^x (y) from both sides, what do you get?

OpenStudy (anonymous):

e^xy'-y'=-e^x(y)

OpenStudy (anonymous):

now subtract e^x to get y'-y' separate?

OpenStudy (anonymous):

No. Then, factorize the left part, what do you get?

OpenStudy (anonymous):

y'(e^x-y')=-e^x(y)

OpenStudy (anonymous):

No. Try again..

OpenStudy (anonymous):

idk what to do with the other y' bc then it will be y'^2

OpenStudy (anonymous):

When you expand y'(e^x-y'), you won't get (e^x y'-y') , so you did the factorization wrong.

OpenStudy (anonymous):

i know i dont know how to do it

OpenStudy (anonymous):

Take out y' as the common, what's the terms left?

OpenStudy (anonymous):

e^x and y'